# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 8. Proof of main theorems
In this section we give proofs to the main theorems of this paper, Theorems 1.4, 1.5, and 1.7. We first prove the quantitative main theorem, Theorem 1.5. Most of the paper is dedicated to this end. Proving it is highly technical and thus for the clarity of the proof we have presented auxiliary results in separate sections, Sections 4, 5, 6, and 7.
After proving Theorem 1.5, we prove the limit form main theorem, Theorem 1.4. It follows straightforwardly from Theorems 1.5 and 8.1.
Proving Theorem 1.7 relies on using Theorems 1.5 and 8.2. The proof is very similar to that of Theorem 1.4 but due to the probabilistic setting imposing lighter conditions on the sequence, we need Theorem 8.2 to record some key results in random surface theory.
We proceed to prove Theorem 1.5. We restate the theorem for the convenience of the reader.
@@
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
@@
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
@@
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
@@
The coefficient appearing is
$C_{8}=50000\pi.$
The proof presented here is clear as the technical details are dealt in different sections, see Sections 6, 5, 6, and 7. The core idea of the proof is to split the problem into a geometric side and a spectral side using a well-chosen propagator. The geometric side and the spectral side can then be dealt separately.
@@
Then due to way the operators are defined, we have that
@@
Using the previous, we can write that
@@
The left factor containing spectral data is bounded using Proposition 4.1,which can be used since $I$ is compact in $(0,\infty)$, $T\delta=\pi/2$, and $\delta$ is smaller than $(2/9)\cdot\min(I)$:
@@
where the Hilbert-Schmidt norm bound is acquired when we consider all the basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
Next we want to apply Proposition 6.1 to bound the Hilbert-Schmidt norm. This requires understanding the integral kernels of $P_{t}$ and $P_{t,\tau}$ for which a convenient form is given in Proposition 5.1. We apply Proposition 6.1 to acquire the following upper bound for the previous expression:
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we opted to use $|X\setminus X(4T)|$ to denote the volume of the points of $X$ with injectivity radius smaller that $4T$. Also,
$C_{7}=\frac{7200}{\pi}.$
Hence we can conclude that
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$C_{7}\cdot 64\pi^{2}\leq 50000\pi=C_{8}.$
We now turn to Theorem 1.4, the limit form main theorem. We also restate the theorem for convenience.
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
@@
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
@@
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
@@
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
@@
Properties (ii) and (iii) remain to be proven; Property (i) was established by Le Masson and Sahlsten (see Theorem 1.1 [LS17]). We remark that the proof presented here for properties (ii) and (iii) could be adapted with slight additions to also include property (i), thus giving a new proof for the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law which is used to asymptotically relate the volume of the surface with the number of eigenvalues in an energy window, when we proceed in the sequence of surfaces. Before the proof, we state the large-scale Weyl law (see [LS17, Sec.9]).
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying property (BSC) stated in Theorem 1.4. $X_{n}$ has Laplace-Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
@@
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
@@
We have included using both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using $\lambda_{j}^{(n)}$ approach. However, the theorem works just as well for $\rho_{j}^{(n)}$ as was seen in the statement of the previous theorem.
@@
Let us define a following function:
@@
$\bar{a}_{n}$ is a mean 0 normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have that
@@
We can apply Theorem 1.5 to the double sum given that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius being smaller than $4T$, $T\,\delta=\pi/2$, and where
@@
The constant appearing is
$C_{8}=50000\pi.$
Then we apply the properties of the sequence to tidy the previous expression. First we note that
@@
And then, since $(a_{n})$ is uniformly bounded – denote this universal bound by $a_{max}$ – we have that
$\|\bar{a}_{n}\|_{\infty}\leq 2\,a_{max}.$
Property (EXP) means that there is a universal lower bound for the spectral gaps which then implies that $\beta(\lambda_{1}^{(n)})$ has a universal lower bound. Denote this by $\beta_{min}$. Finally, Property (UND) means that there is a universal lower bound for the injectivity radii. Denote this by $\operatorname{InjRad}_{min}$. Using these observations gives the following upper bound for Expression (8.2):
$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{8\pi^{2}\,|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{8\pi^{2}\,e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
where the equality followed from splitting the limit supremum into two terms.
Consider the second limit supremum inside the parentheses of the previous expression. Due to the large-scale Weyl law, Theorem 8.1, we have that $N(X_{n},I)$ grows asymptotically like the volume of surfaces $|X_{n}|$. This means that asymptotically the ratio inside the limit supremum closes to 0 as we proceed in the sequence since Property (BSC) forces the ratio of the volume of the points with small injectivity radius and the volume of the surface to go to zero. Hence the previous expression equals
@@
Using the large-scale Weyl law again to the remaining limit supremum, we see that the ratio approaches some constant $C_{9}$. Hence the previous expression equals
$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}$
$=\ \frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
where for the equality we used that $T\,\delta=\pi/2$. If we had chosen originally that
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{8}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\frac{2}{9}\,\min(I)\right),$
we would have that the previous expression is bounded from above by $\epsilon$. We assume that we made this choice in the beginning of the proof to conclude the proof.
In the end we prove the probabilistic theorem, Theorem 1.7. We restate the theorem for convenience.
@@
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
Properties (ii) and (iii) remain to be proven; Property (i) was already proven by Le Masson and Sahlsten [LS24]. The proof is very close to that of Therem 1.4. Now we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced. This is done by the following theorem acquired from the theory of the Weil-Petersson random surfaces.
@@
Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil-Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
@@
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}.$
Then with probability 1o(1)1-o(1) $X$ has the following properties:
@@
$J=\left\{x^{2}+\frac{1}{4}\ :\ x\in I\right\}.$
The theorem is acquired from two articles: [Mon22, HMT25]. Properties (a), (b), and (d) together with the rates is acquired from [Mon22], and Property (c) together with rate is acquired from [HMT25]. When comparing the proof of Theorem 1.7 with the proof of Theorem 1.4, Property (a) replaces Property (UND), Property (b) replaces Property (BSC), and Property (c) replaces Property (EXP). Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten working in this probabilistic setting.
@@
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have that
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
@@
Let us define a following function:
@@
$\bar{a}_{g}$ is a mean 0 normalization of $a_{g}$. Due to orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have that
@@
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)},\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
We can apply Theorem 1.5 to the double sum givem that we assume that $\delta\in(0,(2/9)\cdot\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+8\pi^{2}\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\,\delta=\pi/2$, and where

### v2

@@ 8. Proof of main theorems
In this section, we give proofs of the main theorems of this paper, Theorems 1.3, 1.4, and 1.5. We first prove the quantitative main theorem, Theorem 1.5, as it is needed for the proofs of Theorems 1.3 and 1.4. For clarity of the proof, we have presented auxiliary results in separate sections: Sections 4, 5, 6, and 7.
After proving Theorem 1.5, we prove the limit-form main theorem, Theorem 1.3. This result follows straightforwardly from Theorems 1.5 and 8.1 using Properties (BSC), (EXP), and (UND), where Theorem 8.1 is the large-scale Weyl law.
Proving Theorem 1.4 relies on Theorems 1.5 and 8.2. The proof is very similar to that of Theorem 1.3, but as the probabilistic setting does not impose Properties (BSC), (EXP), and (UND), we need Theorem 8.2 to record some key results in random surface theory to replace these properties.
We proceed to prove Theorem 1.5 and restate the theorem for the convenience of the reader.
@@
Let $X$ be a compact connected hyperbolic surface with Laplace–Beltrami operator $\Delta_{X}$ that has eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values of the eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplication operator. Let
@@
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\,\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, defined by
@@
The coefficient appearing is $C_{8}=5120000\pi^{2}$.
The proof presented here is clear since the technical details are dealt with in different sections; see Sections 4, 5, 6, and 7. The core idea is to split the problem into a geometric side and a spectral side using a well-chosen propagator. These two sides can then be dealt with separately.
@@
Then, due to the way the operators are defined, we have
@@
Using the above, we can write
@@
The left factor containing the spectral data is bounded using Proposition 4.1, which can be applied since $I$ is compact in $(0,\infty)$, $T\delta=\tfrac{\pi}{2}$, and $\delta<\tfrac{2}{9}\min(I)$:
@@
where the Hilbert-Schmidt norm bound is obtained when considering all basis elements. The Hilbert-Schmidt norm contains the geometric data of the problem.
Next, we apply Proposition 6.1 to bound the Hilbert-Schmidt norm:
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{\mathrm{HS}}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ denotes the volume of the points of $X$ with injectivity radius smaller than $4T$. Also, $C_{7}=80000$. Hence we can conclude that
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $C_{7}\cdot 64\pi^{2}=5120000\pi^{2}=C_{8}$.
We now turn to Theorem 1.3, the limit-form main theorem. We also restate the theorem for convenience.
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace–Beltrami operators $(-\Delta_{X_{n}})$. The Laplacian $\Delta_{X_{n}}$ has eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
@@
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplication operator. Then the following properties hold:
@@
    $\lim_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx\right|^{2}=0,$
@@
Properties (ii) and (iii) remain to be proven; Property (i) was established by Le Masson and Sahlsten (see Theorem 1.1 [LS17]). We remark that the proof presented here for Properties (ii) and (iii) could be adapted with slight modifications to also include Property (i), thus giving a new proof of the large-scale quantum ergodicity theorem of Le Masson and Sahlsten.
Passing to the limit and thus acquiring a small upper bound $\epsilon$ is based on carefully using Properties (BSC), (EXP), and (UND) together with the large-scale Weyl law, which asymptotically relates the volume of the surface with the number of eigenvalues in an energy window as we proceed through the sequence of surfaces. Before the proof, we state the large-scale Weyl law (see [LS17, Sec.9]).
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces satisfying Property (BSC) stated in Theorem 1.3. $X_{n}$ has Laplace–Beltrami operator $\Delta_{X_{n}}$ with eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
@@
$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
@@
We include both $\lambda_{j}^{(n)}$ and $\rho_{j}^{(n)}$ since Le Masson and Sahlsten proved their theorem using the $\lambda_{j}^{(n)}$ approach. However, the theorem works just as well for $\rho_{j}^{(n)}$. We especially remark that the large-scale Weyl law follows from the Benjamini–Schramm convergence property, Property (BSC).
@@
Let us define the following function:
@@
$\bar{a}_{n}$ is a mean-zero normalization of $a_{n}$. Due to orthogonality of $\psi_{j}^{(n)}$ and $\psi_{k}^{(n)}$, we have
@@
We can apply Theorem 1.5 to the double sum, given that we assume $\delta\in(0,\tfrac{2}{9}\min(I))$. We get the following upper bound:
$\limsup_{n\to\infty}\ \frac{1}{N(X_{n},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(n)})^{3}}\,\left(\|\bar{a}_{n}\|_{2}^{2}\,+\,\|\bar{a}_{n}\|_{\infty}^{2}\,|X_{n}\setminus X_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right),$ (8.2)
where $|X_{n}\setminus X_{n}(4T)|$ is the volume of the points of $X_{n}$ with injectivity radius smaller than $4T$, $T\delta=\tfrac{\pi}{2}$, and
@@
The constant appearing is $C_{8}=5120000\pi^{2}$.
Then we apply the properties of the sequence to simplify the previous expression. First we note that
@@
Since $(a_{n})$ is uniformly bounded, we have
$\|\bar{a}_{n}\|_{\infty}\leq 2\|a_{n}\|\leq 2\,a_{max},$
where $a_{max}$ is the universal bound for $(a_{n})$. Property (EXP) implies a positive universal lower bound for the spectral gaps, which ensures $\beta(\lambda_{1}^{(n)})$ has a universal positive lower bound. Denote this by $\beta_{min}$. Finally, Property (UND) implies a universal positive lower bound for the injectivity radii. Denote this by $\operatorname{InjRad}_{min}$. Using these observations gives the following upper bound for Expression (8.2):
$\limsup_{n\to\infty}\ 2\cdot C_{8}\,\frac{\max(I)^{4}\,a_{max}^{2}}{N(X_{n},I)\,T\,\beta_{min}^{3}}\left(|X_{n}|+\frac{|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{\operatorname{InjRad}_{min}}\right)$
$=\ \frac{2C_{8}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}\left(\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}\right)+\frac{e^{6T}}{\operatorname{InjRad}_{min}}\,\limsup_{n\to\infty}\left(\frac{|X_{n}\setminus X_{n}(4T)|}{N(X_{n},I)}\right)\right),$
where the equality follows from splitting the limit supremum into two terms.
Consider the second limit supremum inside the parentheses. Due to the large-scale Weyl law, Theorem 8.1, $N(X_{n},I)$ grows asymptotically like the volume $|X_{n}|$. Hence the ratio inside the limit supremum approaches 0 since Property (BSC) forces the volume of points with small injectivity radius relative to the total volume to go to zero. Therefore, the previous expression equals
@@
Using the large-scale Weyl law, the ratio approaches some constant $C_{9}$. Thus we have
$\frac{2C_{8}\,C_{9}\,\max(I)^{4}\,a_{max}^{2}}{T\,\beta_{min}^{3}}=\frac{4\,C_{8}\,C_{9}\,\delta\,\max(I)^{4}}{\beta_{min}^{3}\,\pi},$
where the equality uses $T\delta=\tfrac{\pi}{2}$. If we choose
$\delta=\min\left(\epsilon\cdot\left(\frac{4\,C_{8}\,C_{9}\,\max(I)^{4}}{\beta_{min}^{3}\,\pi}\right)^{-1},\tfrac{2}{9}\,\min(I)\right),$
then the previous expression is bounded by $\epsilon$. We assume this choice in the beginning of the proof to conclude the proof.
In the end, we prove the probabilistic theorem, Theorem 1.4. We restate the theorem for convenience.
@@
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$, and where $g$ is the genus. Let $\Delta_{g}$ be the Laplace–Beltrami operator of $X_{g}$ with eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values of the eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\tfrac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval, and let $(a_{g})$ be a sequence of uniformly bounded functions on $X_{g}$, each acting as a multiplication operator. Then Properties (i)(iii) of Theorem 1.3 hold with high probability as the genus $g$ grows to infinity.
Properties (ii) and (iii) remain to be proven; Property (i) was already proven by Le Masson and Sahlsten [LS24]. The proof is very close to that of Theorem 1.3. Now, we do not assume Properties (BSC), (EXP), and (UND), so these need to be replaced. The following theorem records analogous tools acquired from the theory of Weil–Petersson random surfaces.
@@
Let $X$ be a $\mathbb{P}_{g}$-random surface, where $\mathbb{P}_{g}$ is the Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$ with genus $g$. It has eigenvalues
@@
$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
Then $X$ has the following properties with probability 1o(1)1-o(1) for large genus:
@@
$J=\left\{x^{2}+\tfrac{1}{4}:x\in I\right\}.$
The theorem is taken from two articles: [MON22, HMT25a]. Properties (a), (b), and (d), together with the rates, are taken from [MON22], and Property (c), together with the rate, is taken from [HMT25a]. When comparing the proof of Theorem 1.4 with the proof of Theorem 1.3, Property (a) replaces Property (UND), Property (b) replaces Property (BSC), and Property (c) replaces Property (EXP). Additionally, Property (d) is a suitably modified version of the large-scale Weyl law of Le Masson and Sahlsten, applied in this probabilistic setting.
@@
We prove Properties (ii) and (iii) simultaneously. Let $\epsilon>0$. We aim to find $\delta=\delta(\epsilon)$ such that for every $\tau\in\mathbb{R}$, we have
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)}\in I,\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},a_{g}\psi_{k}^{(g)}\rangle\right|^{2}<\epsilon,$ (8.3)
@@
Let us define the following function:
@@
$\bar{a}_{g}$ is a mean-zero normalization of $a_{g}$. Due to the orthogonality of $\psi_{j}^{(g)}$ and $\psi_{k}^{(g)}$, we have
@@
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\sum_{j:\rho_{j}^{(g)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(g)}\in I,\\ k\neq j,\\ |\rho_{j}^{(g)}-\rho_{k}^{(g)}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j}^{(g)},\bar{a}_{g}\psi_{k}^{(g)}\rangle\right|^{2}.$
We can apply Theorem 1.5 to the double sum, given that we assume $\delta\in(0,\tfrac{2}{9}\min(I))$. We get the following upper bound for the previous expression:
$\limsup_{g\to\infty}\ \frac{1}{N(X_{g},I)}\,C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}^{(g)})^{3}}\,\left(\|\bar{a}_{g}\|_{2}^{2}+\|\bar{a}_{g}\|_{\infty}^{2}\,|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right),$
where $|X_{g}\setminus X_{g}(4T)|$ is the volume of the points of $X_{g}$ with injectivity radius less than $4T$, $T\delta=\tfrac{\pi}{2}$, and where